Calculate the pH of a solution with a [OH-] = 3.0 x 10-6



1
Expert's answer
2021-03-26T05:38:14-0400

Solution:

We can convert between [OH-] and pOH using the following equation:

pOH = - log[OH-]

pOH = - log(3.0×10-6) = 5.52

pOH = 5.52

 

For any aqueous solution at 25C:

pH + pOH = 14

pH = 14 - pOH = 14 - 5.52 = 8.48

pH = 8.48


Answer: pH = 8.48